# HardML: A Benchmark For Evaluating Data Science And Machine Learning knowledge and reasoning in AI

### v1

@@ 2. Data collection
The data collection involved a multi-step process spanned over 6 months. As mentioned in the last paragraph of the introduction, the initial purpose of this project was to form a set of question-answer for ML interview assessment for entrance of the top tech companies. These are to be used on the platform getaiquestions.com, which is a website similar to leetcode.com for interview preparation. Therefore, the dataset construction wasn’t biased towards building problems that the LLMs wouldn’t be able to solve, they were fully intended for human use.
@@ 2.1. The collection pipeline
    Raw data collection and scraping. We have sourced approximately 400 questions from public platforms such as Glassdoor, Blind, Quora, Stack Exchanges, YouTube, as well as from papers and books —including those by Bishop [25]—and from our own writings and public blogs, among many other reputable sources. Note that, we specifically dedicated time in collected ideas from novel and modern sources - very recent interviews on the topic of the latest development in Natural Language Understanding (NLU) or Computer Vision (CV) and collecting ideas from recently published papers (from 2024 and 2023).
@@
    Quality control and data contamination prevention. The is the final step of the process, focusing on rigorous quality assurance and final checks. We meticulously reviewed each multiple-choice question and corresponding answer to ensure accuracy, clarity, and alignment with the benchmark’s objectives. This phase involved collaboration with beta testers—colleagues and peers—who interacted with the questions through the user interface (UI) of our platform (that the project was initially intended for), getaiquestions.com. While no major errors were identified, several ambiguous cases were detected and rectified during this stage, enhancing the robustness and reliability of the final benchmark.
@@ 2.2. Question difficulty
While we acknowledge and we don’t refute that the difficulty assignments may exhibit slight bias—given that they were determined by a single individual—we have strived to represent the difficulties as accurately as possible. This is substantiated by the results of our benchmark evaluations: HardML yields a significantly higher error rate than MMLU, indicating a higher level of difficulty, whereas EasyML achieves a notably lower error rate. These outcomes corroborate our assessments of the relative difficulties of the benchmarks.
@@ 3.3. Sample questions from HardML
Sample problem 1 Question: You want to train a LLM that can solve challenging math problems properly. To do that, you employ a team of mathematicians to devise problems and solutions for training data. Unfortunately, you require a lot of training data, naturally, and hence you have to employ thousands of people to generate problems and solutions for your LLM. You need some form of quality control to understand if the mathematicians keep an overall good quality and that your LLM won’t be trained on corrupted data. You can assume you have 1000 people devising (problem, solution) tasks, one person submits one task. Each task is rated from 5 choices, from 1/5 (lowest) to 5/5 (highest): 1/5,2/5,3/5,4/5,5/5. You want these people to produce, on average, a quality of work of at least 4/5=0.8 and to be 95% sure that is the case. You cannot check all 1000 and compute the average yourself because that would defeat the purpose of employing these people in the first place, so then what’s the minimum number N of random tasks you would need to check? For this exercise, you can assume that the task grades follow a normal distribution and the variance of the overall quality is known and it’s the maximum it can be, given the range 1/5-5/5. Make sure to normalize the grades in [0.2,1] A) 4 B) 6 C) 7 D) 8 Answer: B
@@ 4.1. Accuracy on HardML
Based on a single evaluation of the full benchmark, we found that most models solve aout 65% of the questions with the top performing model (o1) being able to solve 70%. This is in line with other benchmarks from other fields. For example, in math, current benchmarks that are considered ’hard’ like Omni-Math, MathVista and Aime have around 60%, 70% and around 70% respectively in accuracy against o1.
@@ 4.2. Human performance on HardML
    The first metric (Senior Machine Learning Engineer) was obtained during the beta testing phase of data collection (the final step). We invited actual senior machine learning engineers—friends of the authors—to participate in several quizzes consisting of 7 or 8 questions each, sampled from HardML. After aggregating the results, we found that an individual scored, on average, approximately 5.5 correct answers out of 8 (which translates into 68.75%). This performance is reflected in Figure 2. The participants expressed admiration for the benchmark, noting that the questions required significant thought and were highly challenging.
@@ 4.3. Accuracy on MMLU and EasyML
Below, we have the equivalent diagrams for the 112 questions present in the testset of MMLU (ML subset) and our proposed EasyML. Observe how o1 is still the top performer, but the scores are significantly higher compared to HardML. Note that, we have not displayed human assessment figures on the MMLU benchmark as this experiment wasn’t conducted.
@@ 8. Conclusion
With this paper, we instigate to further research in the area of LLM benchmarking for cutting edge Data Science and Machine Learning. The dataset of HardML is present in an interactive environment on getaiquestions.com and can also be obtained in clean json format for experiment replication or further research here. Our work contributes to the ongoing efforts to develop benchmarks that can effectively measure and distinguish the advanced capabilities of AI models in rapidly evolving fields.

### v2

@@ 2. Data collection
The data collection involved a multi-step process spanned over 6 months. As mentioned in the last paragraph of the introduction, the initial purpose of this project was to form a set of question-answer for ML interview assessment for entrance of the top tech companies. These are to be used on the platform neuraprep.com, which is a website similar to leetcode.com for interview preparation. Therefore, the dataset construction wasn’t biased towards building problems that the LLMs wouldn’t be able to solve, they were fully intended for human use.
@@ 2.1. The collection pipeline
    Raw data collection and scraping. We have sourced approximately 400 questions from public platforms such as Glassdoor, Blind, Quora, Stack Exchanges, YouTube, as well as from papers and books —including those by Bishop [11]—and from our own writings and public blogs (ptidor.com), among many other reputable sources. As such, we specifically dedicated time in collected ideas from modern sources - very recent interviews on the topic of the latest development in Natural Language Understanding (NLU) or Computer Vision (CV) and collecting ideas from recently published papers (from 2024 and 2023).
@@
    Quality control and data contamination prevention. The is the final step of the process, focusing on rigorous quality assurance and final checks. We meticulously reviewed each multiple-choice question and corresponding answer to ensure accuracy, clarity, and alignment with the benchmark’s objectives. This phase involved collaboration with beta testers—colleagues and peers—who interacted with the questions through the user interface (UI) of our platform (that the project was initially intended for), neuraprep.com. While no major errors were identified, several ambiguous cases were detected and rectified during this stage, enhancing the robustness and reliability of the final benchmark.
@@ 2.2. Question difficulty
While we acknowledge and don’t refute that the difficulty assignments may exhibit slight bias—given that they were determined by a single individual—we have strived to represent the difficulties as accurately as possible. This is substantiated by the results of our benchmark evaluations: HardML yields a significantly higher error rate than MMLU, indicating a higher level of difficulty, whereas EasyML achieves a notably lower error rate. These outcomes corroborate our assessments of the relative difficulties of the benchmarks. HardML comprises only ”hard” questions (according to the categorization system explained above) whereas MMLU-ML –though lacking an official difficulty rating–appears to consist predominantly of ”easy” and ”medium” questions (according to the same categorization system).
@@ 3.3. Sample questions from HardML
Sample problem 1 Question: You want to train an LLM that can solve challenging math problems properly. To do that, you employ a team of mathematicians to devise problems and solutions for training data. Unfortunately, you require a lot of training data, naturally, and hence you have to employ thousands of people to generate problems and solutions for your LLM. You need some form of quality control to understand if the mathematicians keep an overall good quality and that your LLM won’t be trained on corrupted data. You can assume you have 1000 people devising (problem, solution) tasks, one person submits one task. Each task is rated from 5 choices, from 1/5 (lowest) to 5/5 (highest): 1/5,2/5,3/5,4/5,5/5. You want these people to produce, on average, a quality of work of at least 4/5=0.8 and to be 95% sure that is the case. You cannot check all 1000 and compute the average yourself because that would defeat the purpose of employing these people in the first place, so then what’s the minimum number N of random tasks you would need to check? For this exercise, you can assume that the task grades follow a normal distribution and the variance of the overall quality is known and it’s the maximum it can be (each grade happens at least once), given the range 1/5-5/5. Make sure to normalize the grades in [0.2,1] A) 4 B) 6 C) 7 D) 8 Answer: B
@@ 4.1. Accuracy on HardML
Based on a single evaluation of the full benchmark, we found that most models solve aout 65% of the questions with the top performing model (o1) being able to solve 70%. This is in line with other benchmarks from other fields. For example, in math, current benchmarks that are considered ’hard’ like Omni-Math, MathVista and Aime have around 60%, 70% and around 70% respectively in accuracy against o1. Interestingly, if we allow a ”soft” figure for solved questions (giving partial credit when the model’s answer is a subset of the correct answer), then the performance goes up by 5 percentage points (to 75.08% for o1), not a drastic change.
@@ 4.2. Human performance on HardML
    The first metric (Senior Machine Learning Engineer) was obtained during the beta testing phase of data collection (the final step). We invited actual senior machine learning engineers—friends of the author (see acknowledgements)—to participate in several quizzes consisting of 7 to 8 questions each, sampled, at random, from HardML. Once sampled, the same quizzes were used for each person, we managed to assess 5-6 quizzes per person. After aggregating the results, we found that an individual scored, on average, approximately 5.5 correct answers out of 8 (which translates into 68.75%). Although it is not reflective of the performance on all the questions from HardML (only on a subset), we believe this figure is relevant to be shown. Hence, this performance is reflected in Figure 2 and Figure 3. The participants expressed admiration for the benchmark, noting that the questions required significant thought and were highly challenging.
@@ 4.3. Accuracy on MMLU and EasyML
Below, we have the equivalent diagram (Figure 4) for the 112 questions present in the testset of MMLU (ML subset) and our proposed EasyML. Observe how o1 is still the top performer, but the scores are significantly higher compared to HardML. Note that, we have not displayed human assessment figures on the MMLU benchmark as this experiment wasn’t conducted.
Figure 4. Solved questions in MMLU [ML]
@@ 8. Conclusion
With this paper, we instigate to further research in the area of LLM benchmarking for cutting edge Data Science and Machine Learning. The dataset of HardML is present in an interactive environment on neuraprep.com and can also be obtained in clean json format for experiment replication or further research here. Our work contributes to the ongoing efforts to develop benchmarks that can effectively measure and distinguish the advanced capabilities of AI models in rapidly evolving fields.